OpenStudy (anonymous):

Find the first four terms of the sequence given by the following. a[n]=(-1)^(n+1)・2n^2,n=1,2,3.. HELPPPP ASAPPP

OpenStudy (anonymous):

replace \(n\) by 1 get \[(-1)^{1+1}(1+1)\times (2\times 1^2)\] \[=(-1)^2\times 2\times 2\times 1\] \[=4\]

OpenStudy (anonymous):

others are similar

OpenStudy (campbell_st):

actually the 1st term is \[(-1)^{1 +1} \times 2 \times 1^1\] 2nd term \[(-1)^{2 +1} \times 2 \times 2^2\] and just keep going